STRONG FLUORESCENCE IN A DERIVATIVE OF 2-OXAZOLIN- 5 ONES

\underline{V.B. SINGH}, Department of Physics, Udai Pratap Autonomous College, Varanasi-221002; D.K. RAI, Department of Physics, Banaras Hindu University, Varanasi-221005.

The 2-phenyl-4-(4- methoxy benzylidene) -2-oxazolin-5-one, an important class of hetrocyclic organic molecule, shows strong fluorescence in the ethyl alcohol solution with the excitation of 476.5 nm, 488 nm, 496 nm and 514.5 nm lines of 10 w Ar+ -ion Laser. The photoacoustic spectra and UV-VIS absorption spectra of this organic derivative indicate that the above fluorescence was observed due to singlet-triplet transition. The fluorescence peaks observed at different power and excitation wavelengths indicate the probable laser emission in the above organic derivative.